# Probing Geometrical NSI at the DUNE experiment

### v1

@@ V Summary and Conclusions
The CP discovery potential for a positive value of torsional constant is almost equal to or slightly lower than the standard case excepting for NH-LO and $\delta_{CP}=-90^{o}$ and NH-HO and $\delta_{CP}=+90^{o}$, for which octant-$\delta_{CP}$ degeneracy drives the fit to the wrong octant thus reducing the sensitivity. For negative values of the torsional constant the CP discovery potential is lower than in the standard case.
@@ VI Acknowledgment
Part of the calculations are performed in the High Performance Computing facilities in the SNBNCBS. SG acknowledges the J.C. Bose Fellowship (JCB/2020/000011) by the Anusandhan National Research Foundation, Government of India, and Department of Space Govt. of India. She also acknowledges discussion with Supriya Pan and SNBNCBS for hospitality during the final stage of this work. This research was supported in part by the International Centre for Theoretical Sciences (ICTS) for participating in the programme Understanding the Universe Through Neutrinos (code: ICTS/Neus2024/04).
@@ Appendix A Origin of the interaction
where the sum runs over all species of fermions. The other terms in the action are unchanged, so by varying $\Lambda$ we again get an algebraic equation of motion
@@
$\sum_{i=e,\mu,\tau}\left(\lambda_{i}{\nu}_{iL}^{\dagger}\nu_{iL}\right)\,\tilde{n}\,$. (31)

### v2

@@ IV.4 Octant
In Table 4, we summarize the degeneracies between the four combinations of octant and half-plane of $\delta_{CP}$ in the parameter space, for neutrinos and antineutrinos in the presence of positive, zero, and negative torsion.
Parameters LHP LO LHP HO UHP LO UHP HO
No torsion $\nu$ Degenerate with UHP HO No degeneracy No degeneracy Degenerate with LHP LO
Negative torsion $\nu$ More degeneracy No degeneracy No degeneracy More degeneracy
Positive torsion $\nu$ Less degeneracy No degeneracy No degeneracy Less degeneracy
No torsion $\bar{\nu}$ No degeneracy Degenerate with UHP LO Degenerate with LHP HO No degeneracy
Negative torsion $\bar{\nu}$ No degeneracy Less degeneracy Less degeneracy No degeneracy
Positive torsion $\bar{\nu}$ No degeneracy More degeneracy More degeneracy No degeneracy
Table 4: The octant-$\delta_{CP}$ degenerate parameter space for neutrinos and antineutrinos. Here, LO = Lower octant, HO = Higher octant, UHP = Upper half plane ($0^{\circ}<\delta_{CP}<180^{\circ}$) and LHP = Lower half plane ($180^{\circ}<\delta_{CP}<0^{\circ}$). The positive(negative) torsion takes $\lambda_{(2,3)1}=\pm 0.1\sqrt{G_{F}}$.
@@ V Summary and Conclusions
The CP discovery potential for a positive value of torsional constant is almost equal to or slightly lower than the standard case excepting for NH-LO and $\delta_{CP}=-90^{\circ}$ and NH-HO and $\delta_{CP}=+90^{\circ}$, for which octant-$\delta_{CP}$ degeneracy drives the fit to the wrong octant thus reducing the sensitivity. For negative values of the torsional constant the CP discovery potential is lower than in the standard case.
@@ VI Acknowledgment
Part of the calculations are performed in the High Performance Computing facilities in the SNBNCBS. a. She also acknowledges discussion with Supriya Pan and SNBNCBS for hospitality during the final stage of this work. This research was supported in part by the International Centre for Theoretical Sciences (ICTS) for participating in the programme Understanding the Universe Through Neutrinos (code: ICTS/Neus2024/04).
@@ Appendix A Origin of the interaction
where the sum runs over all species of fermions. More specifically, and since we will be dealing with neutrinos in this work, these fermions should be taken to be in their mass eigenstates, since the Lagrangian of Eq. (23) comes purely from spacetime and is valid in the absence of all other interactions. In principle, one can consider the mass basis to be different from the basis in which the torsional interaction is diagonal, but we will not do that here. The other terms in the action are unchanged, so by varying $\Lambda$ we again get an algebraic equation of motion
@@
$\sum_{i=1,2,3}\left(\lambda_{i}{\nu}_{iL}^{\dagger}\nu_{iL}\right)\,\tilde{n}\,$. (31)